# HP2C-DT: High-Precision High-Performance Computer-enabled Digital Twin

### v1

@@ Abstract
[upcaff]organization=CITCEA, Universitat Politécnica de Catalunya, addressline=Av. Diagonal, 647, Les Corts, city=Barcelona, postcode=08028, state=Catalonia, country=Spain
{highlights}
Introduces HP2C-DT, a reference architecture integrating HPC into digital twins.
Proposes a framework enabling real-time DT execution across edge, cloud, and HPC resources.
Uses COMPSs agents to dynamically orchestrate tasks across heterogeneous infrastructures.
Demonstrates reduced communication overhead while preserving critical information.
Validates fast response times and scalable HPC execution in a power grid use case.
@@ 1 Introduction
A use case of HP2C-DT is in electrical grids, where edge nodes are distributed over vast geographical areas. In such environments, sending data to a central cloud node for analysis is infeasible due to strict latency constraints. Instead, real-time fault detection and response must occur directly at the Edge, while HPC resources handle computationally intensive tasks such as predictive maintenance, load forecasting, and probabilistic risk assessment. Similar scenarios arise in industrial automation, where production line adjustments require millisecond-level reaction times while complex optimization models refine scheduling and resource allocation, and in autonomous transportation networks, where real-time navigation depends on local edge decisions supported by large-scale traffic simulations and data-driven route optimization. Building on the reference architecture and implementation framework, we develop a prototype of a digital twin for power grids to test HP2C-DT’s main features in a real-world scenario.
This article is structured as follows. Section 2 provides a concise review of related work on digital twin architectures, their connection to the computing continuum, and their use of HPC resources. Section 3 outlines the proposed reference software architecture, followed by Section 4, which examines the most important features of the reference architecture in detail. Section 5 introduces the HP2C-DT framework, extending the reference architecture for implementation. Section 6 presents an experimental use case in power systems, demonstrating the benefits of HP2C-DT. Finally, Section 7 provides concluding remarks.
@@ 3 HP2C-DT reference architecture
The Edge layer manages physical devices and local computing resources. It operates autonomously, processing information locally to meet real-time requirements. This layer handles immediate control tasks and real-time monitoring, ensuring quick responses to changes in the environment. For example, sensor measurements collected by edge nodes are transmitted to the Cloud layer using a publish-subscribe messaging model (sensor measurement lines in Figure 1). At the same time, actuation commands from the Cloud are sent back to specific edge nodes through point-to-point messaging (actuation lines in the figure). This dual communication mechanism ensures efficient data collection together with precise control.
@@ 4 Key architectural principles
The HP2C-DT reference architecture balances real-time responsiveness at the Edge with large-scale computational capabilities. This is achieved by integrating operational and computational layers to handle execution, communication, and data flow. This section details the core architectural principles that enable this integration.
@@ 4.2 Data processing management
Functions are modular computation tasks defined by the user in the HP2C-DT architecture. Depending on urgency requirements, they can be configured as synchronous or asynchronous. Because they share the same execution model, they can be used in a similar way at both edge and cloud nodes. Synchronous functions execute locally and immediately on the requesting node. Asynchronous functions, managed by the Parallel and Distributed Execution component, are scheduled and distributed across the digital twin infrastructure in a Function-as-a-Service (FaaS) manner, executing on available resources or offloading to peer nodes. They may consist of a single task or form complex workflows with multiple dependent or nested tasks, each assigned to a node that meets its constraints. In addition to being synchronous or asynchronous, functions can be triggered in different ways depending on system needs:
@@ 4.3 Network traffic reduction
The Edge layer can reduce communication overhead through aggregation methods, which is crucial in high-frequency domains, such as power systems. Each physically bound digital sensor object maintains a rolling window that stores recent measurements. Each entry consists of a timestamp and one or more recorded values, depending on the sensor type.
Rolling windows serve multiple purposes within the digital twin architecture. Functions can leverage them to compute results over recent time frames, but their primary role is to enable aggregation. Without aggregation, sending each measurement separately would overwhelm the network as the number of distributed nodes scales. Instead, edge nodes can transmit entire rolling windows—preserving all data while controlling transmission frequency—or apply aggregation methods that further reduce message size. The latter is particularly useful in environments with a high density of sensor information. In Section 5, we explore the implementations of different aggregation methods.
@@ 5 HP2C-DT framework
The HP2C-DT framework is available for public use [^1] For intellectual property concerns, a link to the repository will be added after the article is accepted for publication.. The following sections go through the most significant features of the framework.
@@ 5.1 COMPSs Agents
The Colony framework [16] introduces an agent-based deployment strategy for COMPSs, allowing any node participating in the infrastructure to automatically convert sequential functions into task-based workflows. By deploying an agent on each node in the infrastructure—whether at the Edge or in the Cloud layers—the HP2C-DT framework facilitates this transformation, allowing functions (described in Section 4.2) to be seamlessly converted into individual COMPSs tasks or orchestrated into structured workflows with appropriate resource allocation. This ensures that code execution fully leverages the available computing resources, including multiple CPU cores, GPUs, and other accelerators embedded in the node. In addition, the framework supports decentralized peer-to-peer orchestration of task executions. Agents collaborate to distribute the workload, allowing a function triggered on the Edge to offload all or part of its workflow to the Cloud layer or neighboring edge nodes. Likewise, since COMPSs is compatible with HPC environments, any task delegated by cloud or edge nodes can be submitted to the HPC system and expanded into a complex, parallel workflow.
@@ 5.2.2 Functions
HP2C-DT functions can be implemented in Java, the architecture’s native language, or Python, which is supported through UNIX domain sockets for local execution. Each function follows a fixed signature, always receiving three arguments: a list or map of sensors it can access, a list or map of actuators it can control, and a map of additional parameters specific to the function’s requirements. This signature allows users to define various computational operations, such as model training, inference, state verification, or other custom computations. Additionally, the implemented triggers (see Section 4.2) allow the execution of functions on demand.
@@
The method-name field provides the fully qualified class path where the logic of the function lies. The parameters section defines the function’s inputs and outputs. In this example, the function processes data from the Voltmeter Gen1 sensor and controls the Three-Phase Switch Gen1 actuator to disconnect the system. Furthermore, an additional parameter, threshold, is listed in other, indicating that the function enforces a voltage limit of 400 volts. The trigger section specifies that this function operates in an event-driven manner, using the onRead mechanism to execute upon receiving new sensor data (refer to Section 4.2 for information about the available function-triggering modes). The trigger-sensor field ensures that only readings from Voltmeter Gen1 initiate execution, while the interval parameter configures the function to activate every five measurements instead of every single reading.
@@ 5.2.3 Aggregation and rolling windows
The framework also provides implementations of aggregation methods, as seen in Section 4.3, specifically the domain-agnostic ones: all, average, sum, and last. The all method is the trivial case where every measurement in the window is sent. sum transmits the sum of all values in the window, average computes and sends the mean, and last retains only the most recent value. Aggregation methods are easily configurable in the edge setup file, as indicated by the aggregate field in Listing 2. In addition, domain-specific aggregation methods may be useful, and the framework’s modular design allows them to be easily integrated.
@@
its phasor representation is $X=Ae^{j\phi}$ or $X=A\angle\phi$. This means that a steady-state sinusoidal signal $x(t)$ can be represented by only two parameters, its amplitude $A$ and its angle $\phi$. We implement the phasor aggregation method using a Discrete Fourier Transform (DFT) approach to process the time series windows.
@@
For last-value aggregation, all previous values are ignored. Assuming an autoregressive process with correlation $\rho$ between consecutive samples, the reconstruction error is:
@@
If $\rho\approx 1$, little information is lost, but for uncorrelated data ($\rho\approx 0$), most information is discarded.
The phasor transformation retains full information for pure sinusoidal signals:
[…]
@@
where $P_{\text{fundamental}}$ is the power of the fundamental frequency component and $P_{\text{total}}$ is the total signal power.
@@ 5.3 Cloud layer implementation
The Cloud layer maintains a collection of virtual digital objects, each representing a physical device linked to an edge node. This collection updates dynamically with every incoming heartbeat and measurement, both of which are managed by the main RabbitMQ broker also running alongside the Cloud. The broker also handles point-to-point actuation towards edge nodes.
@@ 6.1 Case study
This centralized approach has several limitations, including low-bandwidth decision-making, limited automation, and restricted control at the Edge. At the same time, the rise of renewable energy is shifting power systems from a centralized, controllable setup to a more variable and decentralized one. Addressing these limitations requires a more distributed and high-performance approach, which the HP2C-DT framework provides. It supports HPC, which allows for Electromagnetic Transient (EMT) simulations while keeping the traditional power flow and phasor-based models available, ensuring different levels of fidelity. It also enables computations at the Edge, where key assets of power systems—such as generators, power converters, transformers, protection relays, and metering stations—operate as HP2C-DT devices with their corresponding digital replicas.
@@ 6.2 Experimental setup
We use a real-time simulator to model the power grid under different operating scenarios. The OPAL-RT hardware platform, OP4512, [22] runs a real-time EMT model of the electrical network using the Hypersim software [23], capturing dynamics in the kilohertz range. This is essential for accurately modeling power systems dominated by renewable energy. In our prototype, the OPAL-RT hardware represents the actual electrical grid and provides input and output interfaces. We create the necessary digital objects of the implementation-specific layer to interact with it, as explained in Section 5.2.1. Edge nodes run on a computing node directly connected to the OPAL-RT, ensuring real-time interaction. The Cloud layer consists of virtual machines hosted on an on-premise server at BSC, each with 4 CPUs and 8 GB of RAM. HPC resources are provided by the general-purpose partition of the MareNostrum 5 supercomputer, which consists of compute nodes equipped with two Intel Xeon Platinum 8480+ processors—each with 56 cores at 2.0 GHz—for a total of 112 cores per node. Each node includes 512 GB of RAM and is interconnected via NVIDIA HDR InfiniBand for high-throughput, low-latency communication.

### v2

@@ Abstract
[upcaff]organization=CITCEA, Universitat Politècnica de Catalunya, addressline=Av. Diagonal, 647, Les Corts, city=Barcelona, postcode=08028, state=Catalonia, country=Spain
pubnotepubnotefootnotetext: Article accepted in Future Generation Computer Systems. Please cite as:
E. Iraola, M. García-Lorenzo, F. Lordan-Gomis, F. Rossi, E. Prieto-Araujo, R.M. Badia, HP2C-DT: High-Precision High-Performance Computer-enabled Digital Twin, Future Generation Computer Systems, 2025, 108333, ISSN 0167-739X, doi.org/10.1016/j.future.2025.108333.
@@ 1 Introduction
A use case of HP2C-DT is in electrical grids, where edge nodes are distributed over vast geographical areas and operational requirements have evolved considerably with the increasing integration of distributed energy resources interfaced through power electronic converters. Traditional system operation, based on SCADA measurements transmitted to a centralized Transmission System Operator (TSO) control center, was adequate when the grid was dominated by large synchronous generators whose dynamics evolved slowly and were sufficiently captured by multi-second SCADA reporting intervals. The widespread deployment of converter-interfaced resources has introduced faster and more complex transient behaviors that require high-resolution and time-synchronized measurements. Phasor Measurement Units (PMUs), now widely installed within modern Wide Area Monitoring Systems, provide such measurements with reporting rates of  60 samples per second. While these technologies enable real-time observability, real-time detection and mitigation of faults, instabilities, or oscillations require reducing end-to-end latency not only in measurement acquisition but also in decision-making and actuation. For this reason, recent paradigms in power-system monitoring and control advocate for distributed or edge-based computation, in which part of the processing is performed by computational units geographically closer to the measurement devices. By decentralizing computation, the architecture avoids unnecessary communication delays associated with routing all data to the central control room and distributing control signals back to the field, thereby enabling faster local assessment and timely control actions. Within this context, HP2C-DT supports a division of responsibilities in which latency-critical analyses and immediate control actions are performed at the edge, while the centralized computation infrastructure remains responsible for computationally intensive tasks such as predictive maintenance, load and generation forecasting, and probabilistic risk assessment. Similar scenarios arise in industrial automation, where production line adjustments require millisecond-level reaction times while complex optimization models refine scheduling and resource allocation, and in autonomous transportation networks, where real-time navigation depends on local edge decisions supported by large-scale traffic simulations and data-driven route optimization. Building on the reference architecture and implementation framework, we develop a prototype of a digital twin for power grids to test HP2C-DT’s main features in a real-world scenario.
This article is structured as follows. Section 2 provides a concise review of related work on digital twin architectures, their connection to the computing continuum, and their use of HPC resources. Section 3 outlines the proposed reference software architecture, followed by Section 4, which examines the most important features of the reference architecture in detail. Section 5 introduces the HP2C-DT framework, extending the reference architecture for implementation. Section 6 presents an experimental use case in power systems, demonstrating the benefits of HP2C-DT. Section 7 discusses the previous experiments and their results, and examines the scope of the HP2C-DT framework for digital twins within the computing continuum. Finally, Section 8 provides concluding remarks.
@@ 3 HP2C-DT reference architecture
The Edge layer manages physical devices and local computing resources. It operates autonomously, processing information locally to meet real-time requirements. This layer handles immediate control tasks and real-time monitoring, ensuring quick responses to changes in the environment. For example, sensor measurements collected by edge nodes are transmitted to the Cloud layer using a publish-subscribe messaging model (sensor measurement lines in Figure 1). At the same time, actuation commands from the Cloud are sent back to specific edge nodes through point-to-point messaging (actuation lines in the figure).
The decision between publish-subscribe and point-to-point communication arises from the asymmetry in control flow and communication patterns. Edge-to-Cloud communication mainly transmits sensor data upstream, which can be high-frequency, high-volume, and requires many-to-one ingestion. It must also scale with a growing number of edge producers. A publish-subscribe model suits these needs and allows loose coupling between producers and consumers while supporting expansion and modularity. In contrast, Cloud-to-Edge communication carries commands triggered by decisions at the Cloud layer, targeting specific actuations at the Edge. These messages are typically lower in frequency (since higher-frequency actuations are handled locally to enable faster response) and require precise delivery to the correct edge node. This fits a point-to-point communication model, which favors simplicity and determinism, allows each edge node to handle incoming control messages with node-specific logic, and supports bidirectional exchanges through acknowledgments or return responses. While other models such as direct APIs, message queues with explicit routing, or narrowly scoped publish-subscribe topics could technically support these patterns, they often introduce tighter coupling, higher complexity, or scaling overheads that our chosen model avoids in this architecture.
@@ 4 Key architectural principles
The HP2C-DT reference architecture balances real-time responsiveness at the Edge with large-scale computational capabilities. This is achieved by integrating operational and computational layers to handle execution, communication, and data flow. This section details the core architectural principles that enable this integration. Implementation-specific details such as input/output definitions and execution parameters are described in Section 5, where we show how these features are concretely defined and managed in the HP2C-DT framework.
@@ 4.2 Data processing management
Functions are modular computation tasks defined by the user in the HP2C-DT architecture. Depending on urgency requirements, they can be configured as synchronous or asynchronous. Because they share the same execution model, they can be used in a similar way at both edge and cloud nodes. Synchronous functions execute locally and immediately on the requesting node. Asynchronous functions, managed by the Parallel and Distributed Execution component, are scheduled and distributed across the digital twin infrastructure in a Function-as-a-Service (FaaS) manner, executing on available resources or offloading to peer nodes. They may consist of a single task or form complex workflows with multiple dependent or nested tasks, each assigned to a node that meets its constraints.
In addition to being synchronous or asynchronous, functions can be triggered in different ways depending on system needs:
@@
From a state-management perspective, functions in our architecture are designed to be stateless: they compute outputs only from their current inputs. State is encapsulated at the level of Sensor objects (see Section 4.1), which maintain rolling windows or histories of past measurements (see Section 4.3). Functions can be triggered by changes in these sensor states, but they themselves do not persist intermediate data across invocations. This separation of concerns simplifies distribution and scaling, since functions can be executed uniformly across the different architecture layers without managing internal state.
@@ 4.3 Network traffic reduction
The Edge layer can reduce communication overhead through aggregation methods, which is crucial in high-frequency domains, such as power systems. Each physically bound digital sensor object maintains a rolling window that stores recent measurements. Each entry consists of a timestamp and one or more recorded values, depending on the sensor type. The window is updated continuously as new data arrives, discarding the oldest entries to keep a fixed size. This follows the standard sliding-window buffering model commonly used in time-series processing.
Rolling windows serve multiple purposes within the digital twin architecture. Functions can leverage them to compute results over recent time frames, but their primary role is to enable aggregation. Without aggregation, sending each measurement separately would overwhelm the network as the number of distributed nodes scales. Instead, edge nodes can transmit entire rolling windows—preserving all data while controlling transmission frequency—or apply aggregation methods that further reduce message size. The latter is particularly useful in environments with a high density of sensor information. Aggregations may include operations such as averaging, last-value reporting, or domain-specific reductions (e.g., phasor computations). The concrete implementation of these mechanisms is discussed in Section 5.
@@ 5 HP2C-DT framework
The HP2C-DT framework is publicly available at https://github.com/bsc-wdc/HP2C-DT [18]. The following sections go through the most significant features of the framework. It is important to emphasize that the examples used throughout the rest of the article illustrate the framework’s capabilities and potential, but do not represent a comprehensive view of all the functionalities required by a domain-specific, fully operational, end-to-end digital twin.
@@ 5.1 COMPSs Agents
The Colony framework [16] introduces an agent-based deployment strategy for COMPSs, allowing any node participating in the infrastructure to automatically convert sequential functions into task-based workflows. By deploying an agent on each node in the infrastructure—whether at the Edge or in the Cloud layers—the HP2C-DT framework facilitates this transformation, allowing functions (described in Section 4.2) to be seamlessly converted into individual COMPSs tasks or orchestrated into structured workflows with appropriate resource allocation. This ensures that code execution fully leverages the available computing resources, including multiple CPU cores, GPUs, and other accelerators embedded in the node. In addition, the framework supports decentralized peer-to-peer orchestration of task executions. Agents collaborate to distribute the workload, allowing a function triggered on the Edge to offload all or part of its workflow to the Cloud layer or neighboring edge nodes. Each agent does this by checking its own resources to ensure that task scheduling stays within limits, and by evaluating the constraints of each task. For example, some tasks may require a specific processor architecture or a minimum number of CPUs. COMPSs agents handle this automatically.
Likewise, COMPSs agents can be configured with HPC clusters as execution resources, making the HPC layer directly accessible from any node in the framework. This is enabled by a dedicated COMPSs connector for HPC clusters, which interfaces with the underlying batch management systems, handles job submission and tracking—including failure recovery—and manages result serialization and retrieval. Through this mechanism, COMPSs bridges the framework with HPC infrastructures, allowing cloud or edge nodes to offload intensive workloads without additional development effort on the HPC side.
@@ 5.2.2 Functions
HP2C-DT functions can be implemented in Java, the architecture’s native language, or Python, which is supported through UNIX domain sockets for local execution. Python is included here to ease development and give access to scientific and AI libraries that are widely used in digital twin modeling and simulation. Each function follows a fixed signature, always receiving three arguments: a list or map of sensors it can access, a list or map of actuators it can control, and a map of additional parameters specific to the function’s requirements. This signature allows users to define various computational operations, such as model training, inference, state verification, or other custom computations. Additionally, the implemented triggers (see Section 4.2) allow the execution of functions on demand. While HP2C-DT does not assign explicit priorities to triggers, each function is executed in a separate thread, relying on the operating system and JVM scheduler to manage concurrent execution. This approach has proven sufficient for typical workloads, and end users can implement custom prioritization or conflict resolution within their functions if needed.
In terms of interaction with the Parallel and Distributed Execution engine, which in this framework is implemented through COMPSs agents, asynchronous functions are handled in two complementary ways. At the intra-function level, functions are written as ordinary sequential code, and COMPSs transforms them into a workflow based on decorators and synchronization points added by the developer. This remains fully compatible with other parallelization mechanisms used inside the workflow itself, such as BLAS kernels invoked by NumPy’s linear algebra routines or OpenMP parallel regions that can be wrapped as COMPSs tasks. At the higher level, COMPSs schedules the tasks composing the workflows and offloads them across the available nodes. For that, it takes into account both the characteristics of the workload and the hardware resources of each node, which are described in the COMPSs configuration files.
@@
The method-name field provides the fully qualified class path where the logic of the function lies. The parameters section defines the function’s inputs and outputs. In this example, the function processes data from the Voltmeter Gen1 sensor and controls the Three-Phase Switch Gen1 actuator to disconnect the system. Furthermore, an additional parameter, threshold, is listed in other, indicating that the function enforces a voltage limit of 40 kV. The trigger section specifies that this function operates in an event-driven manner, using the onRead mechanism to execute upon receiving new sensor data (refer to Section 4.2 for information about the available function-triggering modes). The trigger-sensor field ensures that only readings from Voltmeter Gen1 initiate execution, while the interval parameter configures the function to activate every five measurements instead of every single reading.
@@ 5.2.3 Aggregation and rolling windows
The framework also provides implementations of aggregation methods, as seen in Section 4.3, specifically the domain-agnostic ones: all, average, sum, and last. The all method is the trivial case where every measurement in the window is sent. sum transmits the sum of all values in the window, average computes and sends the mean, and last retains only the most recent value. We include both sum and average aggregators, which are mathematically closely related, because they are canonical examples of aggregation methods and are commonly supported in practical frameworks. Aggregation methods are easily configurable in the edge setup file, as indicated by the aggregate field in Listing 2. In addition, domain-specific aggregation methods may be useful, and the framework’s modular design allows them to be easily integrated.
@@
its phasor representation is $X=Ae^{j\phi}$ or $X=A\angle\phi$. The frequency is omitted from the phasor representation because a nominal base frequency is assumed to be known and constant, as is standard practice in, e.g., power system analysis. This means that a steady-state sinusoidal signal $x(t)$ with known frequency can be represented by only two parameters, its amplitude $A$ and its angle $\phi$. We implement the phasor aggregation method using a Discrete Fourier Transform (DFT) approach to process the time series windows. The expected frequency of the signal needs to be provided beforehand in the function configuration.
@@
For last-value aggregation, all previous values are ignored. Assuming an autoregressive process with correlation $\rho$ between consecutive samples [20], the reconstruction error is:
@@
The autoregressive correlation $\rho$ is a model of signal smoothness: if consecutive measurements are highly correlated, keeping only the last value retains most of the signal’s information. If $\rho\approx 1$, little information is lost, but for uncorrelated data ($\rho\approx 0$), most information is discarded.
The phasor transformation retains full information for pure sinusoidal signals if the estimated angular frequency, $\omega_{\text{est}}$, aligns with the actual angular frequency, $\omega$}:
$\varepsilon_{\text{phasor}}\approx 0,\>\text{if }x(t)\text{ is a steady-state sinusoid and }\omega_{\text{est}}\approx\omega$ (5)
@@
where $P_{\text{fundamental}}$ is the power of the fundamental frequency component and $P_{\text{total}}$ is the total signal power. This expression measures the fraction of signal energy that the phasor is able to retain when reflecting only the fundamental frequency, which is sufficient for many steady-state applications [19].
@@ 5.3 Cloud layer implementation
The Cloud layer maintains a collection of virtual digital objects, each representing a physical device linked to an edge node. This collection updates dynamically with every incoming heartbeat and measurement, both of which are managed by the main RabbitMQ broker also running alongside the Cloud. The broker also handles point-to-point actuation d
@@ 6.1 Case study
This centralized approach has several limitations, including low-bandwidth decision-making, limited automation, and restricted control at the Edge. At the same time, the rise of renewable energy is shifting power systems from a centralized, controllable setup to a more variable and decentralized one. Addressing these limitations requires a more distributed and high-performance approach, which the HP2C-DT framework provides in the form of a digital twin architecture. It supports the integration of HPC resources and external Electromagnetic Transient (EMT) solvers, enabling EMT simulations while still providing access to traditional power-flow and phasor-based models, thereby ensuring multiple levels of modeling fidelity. It also enables computations at the Edge, where key assets of power systems—such as generators, power converters, transformers, protection relays, and metering stations—operate as HP2C-DT devices with their corresponding digital replicas.
In the domain of power systems, digital twins are implemented as dynamic, real-time replicas of power system assets, supported by connection, data, model, and service layers. Previous efforts in this area typically do not address edge-to-cloud latency limitations, fast reaction, or the integration of HPC resources in the architecture; instead, they emphasize combining machine learning with simulation models for advanced analytics and feedback loops, where physical data refines virtual models, and tackle the challenges of data scarcity, dynamic topology changes, and uncertainties that hinder AI model performance [24].
@@ 6.2 Experimental setup
We use a real-time simulator to model the power grid under different operating scenarios. The OPAL-RT hardware platform, OP4512, [25] runs a real-time EMT model of the electrical network using the Hypersim software [26], capturing dynamics in the kilohertz range. This is essential for accurately modeling power systems dominated by renewable energy. In our prototype, the OPAL-RT hardware represents the actual electrical grid and provides input and output interfaces. We create the necessary digital objects of the implementation-specific layer to interact with it, as explained in Section 5.2.1.
Edge nodes run on a workstation directly connected to the OPAL-RT, ensuring real-time interaction. The workstation runs Ubuntu 22.04 and is equipped with an Intel i7-1265U CPU and 16 GB of RAM. Communication with the Cloud layer takes place over the institutional network using standard TCP/IP, with RabbitMQ as the messaging middleware. The workstation connects to the local network via Wi-Fi. The Cloud layer consists of virtual machines hosted on an on-premise server at BSC, each with 4 CPUs and 8 GB of RAM. Measured average round-trip latency between edge and cloud nodes is approximately 30 ms. HPC resources are provided by the general-purpose partition of the MareNostrum 5 supercomputer, which consists of compute nodes equipped with two Intel Xeon Platinum 8480+ processors—each with 56 cores at 2.0 GHz—for a total of 112 cores per node. Each node includes 512 GB of RAM and is interconnected via NVIDIA HDR InfiniBand for high-throughput, low-latency communication.